# Understanding the Challenges of Maker Entrepreneurship

### v1

@@ 2.2.2. Entrepreneurship in HCI literature
For instance, many have highlighted the importance of offline social support when acquiring entrepreneurial and technology skills (Casey, 2014; Hui et al., 2020; Dillahunt et al., 2022; Kotturi et al., 2022; Hui and Gerber, 2017; Lindtner et al., 2014). Forming offline social support groups can make it easier for entrepreneurs to learn how their peers use technology for their business and exchange trustworthy advice (Hui et al., 2020). More specifically, offline social support offers intention and authenticity among entrepreneurs, which can make it easier to vet peers as compared to online interactions  (Kuhn et al., 2016). However, it remains unclear whether and how the challenges of digital entrepreneurial tools apply to maker entrepreneurs in particular. Therefore, we extend the research in HCI on entrepreneurship by detailing how maker entrepreneurs use digital tools and engage in online and offline social networks to provide support for each others’ tool discovery and tool use.
@@ 2.3. Research gap: business support technologies for maker entrepreneurs
“Digital technologies play a central role in bridging access to resources, connections, and opportunity” (pg. 1) (Avle et al., 2017) Entrepreneurship is heavily mediated by technology, thus CSCW research (Annett, 2020; Kotturi et al., 2024) has been concerned with understanding who these people are and learning how they achieve their goals. Many of the work has focused on people in rural and under-resourced environments. Much of the work in CSCW and HCI has focused on understanding their populations, as well as a few works that developed new tools (Kotturi et al., 2024). Our work focuses on creative entrepreneurship, since this is lightly discussed. Creative entrepreneurship has similarities to micro-entrepreneurship, but we are emphasizing the differences.
Tools that are designed to help entrepreneurs have been critiqued to disproportionately favor people who have entrepreneurial advantages (Dillahunt et al., 2018). To date, prior technology research in the creative entrepreneurship space has overwhelmingly focused on creativity support tools for those engaged in open-ended creative work, but this research does not cover creative support tools with business-oriented functions (i.e., finances, operations, marketing). Instead, the creative support tools have features like brainstorming  (Foong et al., 2017), getting feedback on creations (Kotturi et al., 2021, 2024), self-presentation (Crain and Bailey, 2017; Xu et al., 2015; Dabbish et al., 2012; Kim et al., 2017), and setting open-ended goals (Krishna Kumaran et al., 2021). Research on creativity tools often does not touch on the business-oriented functions where makers may have the least training and knowledge. This poses a challenge for budding maker entrepreneurs, who often work in isolation of supportive and knowledge-sharing communities and have difficulties identifying and remedying gaps in their business knowledge (Shultz, 2015).
@@ 4.1.1. Reconciling maker and entrepreneur identities
A recurring theme in our interviews was the tension maker entrepreneurs felt between making and entrepreneurship and the values that they attempted to prioritize. For some, entrepreneurship was a means to the end of making—5 out of 20 makers tended to explicitly look for ways to prioritize or give more time to their craft, rather than maximize financial reward. P22, a potter, said, “I wanted to sell my art first and I did it. I wanted to get the artist’s residency and I got it. I wanted to put my pieces in a gallery, which I got it. But my goal now is to put my pieces in a museum in the future. That is my 10-year goal.” For P22, the financial gain of selling their work allows them to work for their higher artistic goals. Other interviewees were quick to proclaim other priorities they put ahead of money. For instance, one of our consultants, who runs a business incubator, I06, said: I struggle with that phrase, [entrepreneurship]. I think I’m an anti-capitalist in my heart of hearts. I don’t have any interest in being a millionaire or in helping other people to become millionaires… Like, I think that there are a lot of reasons that people would not want to scale a business or make millions with a business, and that doesn’t mean that they shouldn’t get support for it. In this, I06 recognizes that many maker entrepreneurs may want to run a stable business, but may not have goals to grow large, suggesting maker enterpeneurs with different values may need sufficient, yet a different kind of support.
While turning a profit and running a sustainable business need not be in opposition to creative maker work, makers faced challenges combining the maker and business identity, psychologically complicating the enjoyment of creative activity. P07, a lamp maker, stated: “I’m trying to figure out, like, do I make products to sell to people or do I just make things for my own kind of happiness and fulfillment?” For some, such as P11, a stationery maker, this tension ultimately led to their reverting to making as a hobby rather than a business venture: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] So I’ve kind of since turned that into more of a hobby…rather than business…That kind of took away the joy of what I was trying to do with it, and it just like became too stressful, too much pressure—too much expense upfront, with very little knowing that I could actually make that money back. These quotes suggest the tension makers may feel when figuring out how to balance their creative goals and enjoyment with the need to run a sustainable business.
@@ 4.1.3. Grappling with business basics
Our interviewees’ difficulties reconciling their values and priorities in entrepreneurship, and their ”accidental” entrance into entrepreneurship, also seemed to affect their business operations. By and large, the maker entrepreneurs in our study did not keep consistent records of their receipts, production costs, and taxes at the early stages of their business due to a “side gig mindset.” A recurring theme in our results was maker entrepreneurs managing the business side of the work in an ad-hoc manner. For example, P07 tracked payments from customers through multiple platforms which he manually updated in an application not originally intended to support business transactions: “they just Venmo me… and then I keep their names in a [iPhone] Notes app.” The entrepreneurial hub co-founder, I04, with finance expertise, corroborated that accounting was a particular shortcoming with many creative entrepreneurs: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] Nobody thinks about accounts receivable, taxes, and all of these things that really matter… it’s beneficial to think about those things starting out, instead of only thinking about the logo…
@@
{adjustwidth}
20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] When I first started kind of just making money, it was just strictly cash. I didn’t know how to write invoices. It was just, pull up to the shoot with the cash ready… I learned over time that the proper way to do business is to learn how to write invoices.
@@ 4.2.1. Challenges of small-scale physical production
However, small-scale physical production gave the maker entrepreneurs we spoke to a sense of fulfillment and a satisfying connection to the made artifact, again speaking to the values they may have over pure financial growth. P25, a gem cutter, describes that tangible creations give her satisfaction outside of her slower-paced, full-time job as a scientist, where results took time: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] You know, gem cutting, I can sit down in an evening, finish a gem, hand it off to a satisfied customer, you know, they come back in a few months and show me what they’ve made and it’s very fulfilling to see from the beginning to the end and you see you have a tangible impact on someone’s life. Thus, changing production processes may not be a desired option for makers as they grow their business.
@@ 4.2.2. Connecting to clients
However, the right client who finds the makers work could be transformative for the maker’s business. P29, a maker who turned vintage tennis rackets into guitars, described one such case: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] I’m hoping that I can build up a following; where people go, why the hell did you make this? Or, holy shit, I want one! …you’re not trying to sell to everyone immediately. …I’m gonna give one to my artist friend who’s gonna make a bunch of content with it. So he’s the guy that’s gonna make a song out of it. He’s the guy. He has a big following. Other maker entrepreneurs “shopped” for clients who could afford their products to support a sustainable business model. P12, for example, described trying to attract wealthier clients: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] Our goal is to be able to attract [wedding] clients that are willing to pay us more. We know that those clients are also willing to invest more into the [photography] open-ended planning process and that will influence the quality of our work. Similarly, the gift basket maker, P13, customized her baskets based on how much customers could pay. Ultimately, these varying strategies suggest that maker entrepreneurs took different strategies in finding customers over just connecting with people admiring their work.
@@ 4.2.3. Pricing their goods
In addition to finding paying customers, entrepreneurs also had a variety of strategies for pricing their goods; they looked at the perceived value in the market, the size of the piece, client spending power, and the cost of materials and sourcing location. The makers we spoke with, however, had unique challenges because craft goods are somewhere between the spectrum between commodity goods, which are priced based on market value, and art, which is priced based on the symbolic value (Khaire, 2017)—far more elusive to pin down.
@@
We saw some indications that successful maker entrepreneurs we spoke with were able to reason about and capitalize on the non-monetary value that customers derive from maker goods. The painter P02 discussed the values associated with nostalgia: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] You know, they’ll play it on a cassette, and they’ll be nostalgic, and maybe they’ll buy more cassettes. So it’s like, people buy this for \$15 right now [and] they keep this. This is like a low-priced item. They keep it for 10 years. And as I’m working and doing my thing, and it’s like, oh, shoot, like you had his first mixtape. P02 understood that items he made could accumulate value over time because of nostalgia.
P09 also described the way that consumers purchased items as an expression of their values: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] But like Tom’s [^1] Tom’s Shoes is a company that donates a pair of shoes for each pair sold.Giving away shoes if you buy shoes like really makes people feel good…Impacting somebody instead of, you know, Amazon or Apple. And then I think, yeah, we express ourselves through our consumption. Building on this idea of thinking about pricing with regard to how customers express themselves, I02, who ran hands-on workshops for various maker activities such as woodworking, knife-making, and jewelry, described his value to his customers as selling them “bragging rights”: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] They get to experience what it’s like to weld even though they’re a marketing professional who dresses nicely and goes to respectable places. They come here and get to make a lot of noise and sparks… hopefully their friends notice [the thing they made], and they get to brag on themselves and say, I made that. Through this experience, customers got to tell their friends about their new skills as well as develop a new sensibility about the products they bought. Based on these comments from our interviewees, there is an opportunity for maker entrepreneurs to develop the sophistication to reason about the non-monetary value that customers derive from their goods. With such sophistication, maker entrepreneurs can better design products that engage with these values, such as the nostalgia, bragging rights, and self-expression mentioned above.
@@ 4.2.4. Building a community
For maker entrepreneurs, community and relationships were interwoven across their personal identity and their business brand. Building a community of friends and followers who admire their work on social media doubles as a potential customer base and enables them to create close relationships with customers, suppliers, and business partnerships. Our interviewees recognized the importance of community and discussed how they created community (P02, P04, P06, P09, P11, P12, P20, I04).
@@
Participants reflected on these kinds of relationships which required balancing genuine relationships that were personally fulfilling while also ensuring business success. One curator, I07, we spoke with described the ideal: “I wanted to promote artists and wanted to have a great…connection and relationship with artist[s]” suggesting how positive connections could lead to more business through brokered promotion. However, for some entrepreneurs, relationships could interfere with a more objective assessment of their products. As P11 described: “I just wanted to test [my idea] out without having my name seeded with it so that people wouldn’t feel like, ‘I’ll do this to support [P11’s name].”’ Here, P11 described how they struggled to garner honest feedback from their community that was not cushioned with friendly support, making it harder for them to test if their idea had true market viability. Such signals of customer desire were critical for maker entrepreneurs, who, due to their small-scale, had limited time to dedicate to research and development. Our findings suggest that systems to support maker entrepreneurs through community should be designed to assist creating connections that support business, beyond simply making connections between people and the maker.
@@ 4.3.2. Finding customers
However, while makers were aware of the algorithms that could promote or bury their posts and regulate communication with their fans and customers, they were not always keen to engage with specific posting and promotion practices. Some maker entrepreneurs described using platforms that were less algorithmic, so that they could have more control over their communication with customers and fans. P25 described how they used Reddit to promote their work because the Instagram algorithm could change and reduce the number of people who see their work. This in turn could lead to a reduction in sales that the maker had little to no control over, since they did not understand how the algorithm had changed. They also described how the design of commenting on Instagram could limit engagement: {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] [The Reddit] community is so much more of a draw to me than just posting on Instagram by myself… I really like being able to better interact with clients and have that platform. In Instagram comments, if you write a comment, you’re just scrolling, you see like two comments, and it’s really hard to follow threads. P25’s challenges with Instagram were echoed by P07, who described Instagram as primarily being a one-way communication tool for marketing and lamenting that they felt there were no great platforms for two-way communication between maker entrepreneurs and their customers. For these maker entrepreneurs though, such communication was important for finding customers and marketing their work.
@@ 4.3.4. Managing finances & operations
While all businesses that sell physical goods need to track their finances, inventory, and communications, many of our participants found technology solutions lacking for specific aspects of their maker business. Regarding financial tracking, we found that participants using more formal sales tools used the financial tracking that came with them. Some participants mentioned using professional tools such as Quicken [^2] https://www.quicken.com (P23) and Manager [^3] https://www.manager.io (P19) for tracking financial information, and one participant managed their own financial spreadsheet; however, these were exceptions among our interviewees. P27, the hair-clip designer with a background in economics, was an anomaly, as she enjoyed doing taxes and keeping her receipts because it gave her control of her business. She recommended that all entrepreneurs should keep a shoe-box [^4] Storing receipts in a shoe-box is a common strategy among personal finance recommendations. with their receipts. She explained, {adjustwidth}20pt20pt {mdframed}[linecolor=black, linewidth=1pt, topline=false, bottomline=false, rightline=false, leftline=true, backgroundcolor=white, innerrightmargin=0pt, ] Quicken makes something where you can just take a picture….It’s all on your phone and [categorizes] what kind of expense it is…But I think if you don’t do it yourself, you don’t know what receipts to save. She alludes to an important point here - that an efficient tool does not necessarily create financial knowledge. Instead, financial knowledge supports the use of efficient tools.

### v2

@@ 2.2.2. Entrepreneurship in HCI literature
For instance, many have highlighted the importance of offline social support when acquiring entrepreneurial and technology skills (Casey, 2014; Hui et al., 2020; Dillahunt et al., 2022; Kotturi et al., 2022; Hui and Gerber, 2017; Lindtner et al., 2014). Forming offline social support groups can make it easier for entrepreneurs to learn how their peers use technology for their business and exchange trustworthy advice (Hui et al., 2020). More specifically, offline social support offers intention and authenticity among entrepreneurs, which can make it easier to vet peers as compared to online interactions  (Kuhn et al., 2016). However, it remains unclear whether and how the challenges of digital entrepreneurial tools apply to maker entrepreneurs in particular.
Therefore, we extend the research in HCI on entrepreneurship by detailing how maker entrepreneurs use digital tools and engage in online and offline social networks to provide support for each others’ tool discovery and tool use.
@@ 2.3. Research gap: business support technologies for maker entrepreneurs
“Digital technologies play a central role in bridging access to resources, connections, and opportunity” (pg. 1) (Avle et al., 2017) Entrepreneurship is heavily mediated by technology, thus CSCW research (Annett, 2020; Kotturi et al., 2024) has been concerned with understanding who these people are and learning how they achieve their goals. Many of the work has focused on people in rural and under-resourced environments.
Much of the work in CSCW and HCI has focused on understanding their populations, as well as a few works that developed new tools (Kotturi et al., 2024). Our work focuses on creative entrepreneurship, since this is lightly discussed. Creative entrepreneurship has similarities to micro-entrepreneurship, but we are emphasizing the differences.
Tools that are designed to help entrepreneurs have been critiqued to disproportionately favor people who have entrepreneurial advantages (Dillahunt et al., 2018). To date, prior technology research in the creative entrepreneurship space has overwhelmingly focused on creativity support tools for those engaged in open-ended creative work, but this research does not cover creative support tools with business-oriented functions (i.e., finances, operations, marketing). Instead, the creative support tools have features like brainstorming  (Foong et al., 2017), getting feedback on creations (Kotturi et al., 2021, 2024), self-presentation (Crain and Bailey, 2017; Xu et al., 2015; Dabbish et al., 2012; Kim et al., 2017), and setting open-ended goals (Krishna Kumaran et al., 2021).
Research on creativity tools often does not touch on the business-oriented functions where makers may have the least training and knowledge. This poses a challenge for budding maker entrepreneurs, who often work in isolation of supportive and knowledge-sharing communities and have difficulties identifying and remedying gaps in their business knowledge (Shultz, 2015).
@@ 4.1.1. Reconciling maker and entrepreneur identities
A recurring theme in our interviews was the tension maker entrepreneurs felt between making and entrepreneurship and the values that they attempted to prioritize. For some, entrepreneurship was a means to the end of making—5 out of 20 makers tended to explicitly look for ways to prioritize or give more time to their craft, rather than maximize financial reward. P22, a potter, said, “I wanted to sell my art first and I did it. I wanted to get the artist’s residency and I got it. I wanted to put my pieces in a gallery, which I got it. But my goal now is to put my pieces in a museum in the future. That is my 10-year goal.” For P22, the financial gain of selling their work allows them to work for their higher artistic goals.
Other interviewees were quick to proclaim other priorities they put ahead of money. For instance, one of our consultants, who runs a business incubator, I06, said: I struggle with that phrase, [entrepreneurship]. I think I’m an anti-capitalist in my heart of hearts. I don’t have any interest in being a millionaire or in helping other people to become millionaires… Like, I think that there are a lot of reasons that people would not want to scale a business or make millions with a business, and that doesn’t mean that they shouldn’t get support for it. In this, I06 recognizes that many maker entrepreneurs may want to run a stable business, but may not have goals to grow large, suggesting maker enterpeneurs with different values may need sufficient, yet a different kind of support.
While turning a profit and running a sustainable business need not be in opposition to creative maker work, makers faced challenges combining the maker and business identity, psychologically complicating the enjoyment of creative activity. P07, a lamp maker, stated: “I’m trying to figure out, like, do I make products to sell to people or do I just make things for my own kind of happiness and fulfillment?” For some, such as P11, a stationery maker, this tension ultimately led to their reverting to making as a hobby rather than a business venture: So I’ve kind of since turned that into more of a hobby…rather than business…That kind of took away the joy of what I was trying to do with it, and it just like became too stressful, too much pressure—too much expense upfront, with very little knowing that I could actually make that money back. These quotes suggest the tension makers may feel when figuring out how to balance their creative goals and enjoyment with the need to run a sustainable business.
@@ 4.1.3. Grappling with business basics
Our interviewees’ difficulties reconciling their values and priorities in entrepreneurship, and their ”accidental” entrance into entrepreneurship, also seemed to affect their business operations. By and large, the maker entrepreneurs in our study did not keep consistent records of their receipts, production costs, and taxes at the early stages of their business due to a “side gig mindset.” A recurring theme in our results was maker entrepreneurs managing the business side of the work in an ad-hoc manner. For example, P07 tracked payments from customers through multiple platforms which he manually updated in an application not originally intended to support business transactions: “they just Venmo me… and then I keep their names in a [iPhone] Notes app.” The entrepreneurial hub co-founder, I04, with finance expertise, corroborated that accounting was a particular shortcoming with many creative entrepreneurs: Nobody thinks about accounts receivable, taxes, and all of these things that really matter… it’s beneficial to think about those things starting out, instead of only thinking about the logo…
@@
When I first started kind of just making money, it was just strictly cash. I didn’t know how to write invoices. It was just, pull up to the shoot with the cash ready… I learned over time that the proper way to do business is to learn how to write invoices.
@@ 4.2.1. Challenges of small-scale physical production
However, small-scale physical production gave the maker entrepreneurs we spoke to a sense of fulfillment and a satisfying connection to the made artifact, again speaking to the values they may have over pure financial growth. P25, a gem cutter, describes that tangible creations give her satisfaction outside of her slower-paced, full-time job as a scientist, where results took time: You know, gem cutting, I can sit down in an evening, finish a gem, hand it off to a satisfied customer, you know, they come back in a few months and show me what they’ve made and it’s very fulfilling to see from the beginning to the end and you see you have a tangible impact on someone’s life. Thus, changing production processes may not be a desired option for makers as they grow their business.
@@ 4.2.2. Connecting to clients
However, the right client who finds the makers work could be transformative for the maker’s business. P29, a maker who turned vintage tennis rackets into guitars, described one such case: I’m hoping that I can build up a following; where people go, why the hell did you make this? Or, holy shit, I want one! …you’re not trying to sell to everyone immediately. …I’m gonna give one to my artist friend who’s gonna make a bunch of content with it. So he’s the guy that’s gonna make a song out of it. He’s the guy. He has a big following.
Other maker entrepreneurs “shopped” for clients who could afford their products to support a sustainable business model. P12, for example, described trying to attract wealthier clients: Our goal is to be able to attract [wedding] clients that are willing to pay us more. We know that those clients are also willing to invest more into the [photography] open-ended planning process and that will influence the quality of our work. Similarly, the gift basket maker, P13, customized her baskets based on how much customers could pay. Ultimately, these varying strategies suggest that maker entrepreneurs took different strategies in finding customers over just connecting with people admiring their work.
@@ 4.2.3. Pricing their goods
In addition to finding paying customers, entrepreneurs also had a variety of strategies for pricing their goods; they looked at the perceived value in the market, the size of the piece, client spending power, and the cost of materials and sourcing location.
The makers we spoke with, however, had unique challenges because craft goods are somewhere between the spectrum between commodity goods, which are priced based on market value, and art, which is priced based on the symbolic value (Khaire, 2017)—far more elusive to pin down.
@@
We saw some indications that successful maker entrepreneurs we spoke with were able to reason about and capitalize on the non-monetary value that customers derive from maker goods. The painter P02 discussed the values associated with nostalgia: You know, they’ll play it on a cassette, and they’ll be nostalgic, and maybe they’ll buy more cassettes. So it’s like, people buy this for \$15 right now [and] they keep this. This is like a low-priced item. They keep it for 10 years. And as I’m working and doing my thing, and it’s like, oh, shoot, like you had his first mixtape. P02 understood that items he made could accumulate value over time because of nostalgia.
P09 also described the way that consumers purchased items as an expression of their values: But like Tom’s [^1] Tom’s Shoes is a company that donates a pair of shoes for each pair sold.Giving away shoes if you buy shoes like really makes people feel good…Impacting somebody instead of, you know, Amazon or Apple. And then I think, yeah, we express ourselves through our consumption.
Building on this idea of thinking about pricing with regard to how customers express themselves, I02, who ran hands-on workshops for various maker activities such as woodworking, knife-making, and jewelry, described his value to his customers as selling them “bragging rights”: They get to experience what it’s like to weld even though they’re a marketing professional who dresses nicely and goes to respectable places. They come here and get to make a lot of noise and sparks… hopefully their friends notice [the thing they made], and they get to brag on themselves and say, I made that. Through this experience, customers got to tell their friends about their new skills as well as develop a new sensibility about the products they bought. Based on these comments from our interviewees, there is an opportunity for maker entrepreneurs to develop the sophistication to reason about the non-monetary value that customers derive from their goods. With such sophistication, maker entrepreneurs can better design products that engage with these values, such as the nostalgia, bragging rights, and self-expression mentioned above.
@@ 4.2.4. Building a community
For maker entrepreneurs, community and relationships were interwoven across their personal identity and their business brand. Building a community of friends and followers who admire their work on social media doubles as a potential customer base and enables them to create close relationships with customers, suppliers, and business partnerships.
Our interviewees recognized the importance of community and discussed how they created community (P02, P04, P06, P09, P11, P12, P20, I04).
@@
Participants reflected on these kinds of relationships which required balancing genuine relationships that were personally fulfilling while also ensuring business success. One curator, I07, we spoke with described the ideal: “I wanted to promote artists and wanted to have a great…connection and relationship with artist[s]” suggesting how positive connections could lead to more business through brokered promotion.
However, for some entrepreneurs, relationships could interfere with a more objective assessment of their products. As P11 described: “I just wanted to test [my idea] out without having my name seeded with it so that people wouldn’t feel like, ‘I’ll do this to support [P11’s name].”’ Here, P11 described how they struggled to garner honest feedback from their community that was not cushioned with friendly support, making it harder for them to test if their idea had true market viability. Such signals of customer desire were critical for maker entrepreneurs, who, due to their small-scale, had limited time to dedicate to research and development. Our findings suggest that systems to support maker entrepreneurs through community should be designed to assist creating connections that support business, beyond simply making connections between people and the maker.
@@ 4.3.2. Finding customers
However, while makers were aware of the algorithms that could promote or bury their posts and regulate communication with their fans and customers, they were not always keen to engage with specific posting and promotion practices. Some maker entrepreneurs described using platforms that were less algorithmic, so that they could have more control over their communication with customers and fans. P25 described how they used Reddit to promote their work because the Instagram algorithm could change and reduce the number of people who see their work. This in turn could lead to a reduction in sales that the maker had little to no control over, since they did not understand how the algorithm had changed. They also described how the design of commenting on Instagram could limit engagement: [The Reddit] community is so much more of a draw to me than just posting on Instagram by myself… I really like being able to better interact with clients and have that platform. In Instagram comments, if you write a comment, you’re just scrolling, you see like two comments, and it’s really hard to follow threads. P25’s challenges with Instagram were echoed by P07, who described Instagram as primarily being a one-way communication tool for marketing and lamenting that they felt there were no great platforms for two-way communication between maker entrepreneurs and their customers. For these maker entrepreneurs though, such communication was important for finding customers and marketing their work.
@@ 4.3.4. Managing finances & operations
While all businesses that sell physical goods need to track their finances, inventory, and communications, many of our participants found technology solutions lacking for specific aspects of their maker business. Regarding financial tracking, we found that participants using more formal sales tools used the financial tracking that came with them. Some participants mentioned using professional tools such as Quicken [^2] https://www.quicken.com (P23) and Manager [^3] https://www.manager.io (P19) for tracking financial information, and one participant managed their own financial spreadsheet; however, these were exceptions among our interviewees. P27, the hair-clip designer with a background in economics, was an anomaly, as she enjoyed doing taxes and keeping her receipts because it gave her control of her business. She recommended that all entrepreneurs should keep a shoe-box [^4] Storing receipts in a shoe-box is a common strategy among personal finance recommendations. with their receipts. She explained, Quicken makes something where you can just take a picture….It’s all on your phone and [categorizes] what kind of expense it is…But I think if you don’t do it yourself, you don’t know what receipts to save. She alludes to an important point here - that an efficient tool does not necessarily create financial knowledge. Instead, financial knowledge supports the use of efficient tools.